# Speech-Language Models with Decoupled Tokenizers and Multi-Token Prediction

### v1

@@ Abstract
Speech-language models (SLMs) offer a promising path toward unifying speech and text understanding and generation. However, challenges remain in achieving effective cross-modal alignment and high-quality speech generation. In this work, we systematically investigate the impact of key components (i.e., speech tokenizers, speech heads, and speaker modeling) on the performance of LLM-centric SLMs. We compare coupled, semi-decoupled, and fully decoupled speech tokenizers under a fair SLM framework and find that decoupled tokenization significantly improves alignment and synthesis quality. To address the information density mismatch between speech and text, we introduce multi-token prediction (MTP) into SLMs, enabling each hidden state to decode multiple speech tokens. This leads to up to 12 faster decoding and a substantial drop in word error rate (from 6.07 to 3.01). Furthermore, we propose a speaker-aware generation paradigm and introduce RoleTriviaQA, a large-scale role-playing knowledge QA benchmark with diverse speaker identities. Experiments demonstrate that our methods enhance both knowledge understanding and speaker consistency. 111Our demo, code, models, and the RoleTriviaQA dataset will be available at https://github.com/cnxupupup/SLM-Decoupled-MTP.
Speech-Language Models with Decoupled Tokenizers and
Multi-Token Prediction
Xiaoran Fan1††thanks: Equal contribution., Zhichao Sun111footnotemark: 1, Yangfan Gao111footnotemark: 1, Jingfei Xiong111footnotemark: 1, Hang Yan211footnotemark: 1, Yifei Cao1, Jiajun Sun1, Shuo Li1, Zhihao Zhang1, Zhiheng Xi1, Yuhao Zhou1, Senjie Jin1, Changhao Jiang1, Junjie Ye1, Ming Zhang1, Rui Zheng1, Zhenhua Han, Yunke Zhang3, Demei Yan3, Shaokang Dong3, Tao Ji1††thanks: Corresponding author., Tao Gui122footnotemark: 2, Qi Zhang122footnotemark: 2, Xuanjing Huang122footnotemark: 2 1Fudan University 2The Chinese University of Hong Kong 3Honor Device Co., Ltd {xrfan23, taoji, tgui, qz, xjhuang}@fudan.edu.cn
@@ 1 Introduction
In recent years, large language models (LLMs) have achieved remarkable progress in text understanding and generation Grattafiori et al. (2024); Qwen et al. (2025), fueling rapid advancements in multimodal models such as vision-language models Liu et al. (2023), vision-language-action models Kim et al. (2024), and speech-language models (SLMs) Chu et al. (2023); Li et al. (2025). Among them, SLMs are particularly valuable for applications like natural and fluent human-computer dialogue Défossez et al. (2024); Xu et al. (2025) and personalized speech generation Du et al. (2024a, c). The core challenge lies in generalizing the language understanding and generation capabilities of LLMs to incorporate and integrate speech understanding and generation.
Current approaches for connecting text and speech in SLMs can be broadly categorized into two perspectives. Speech-centric models are built upon mature speech generation architectures, with LLMs introduced as conditional context. However, while this approach achieves high-quality speech synthesis, it struggles to fully leverage LLMs’ rich world knowledge and powerful capabilities. LLM-centric models take LLMs as the backbone and extend them with speech interaction. In theory, it fully utilizes LLMs’ strengths (e.g., intent understanding, instruction following, planning, and decision-making). However, the quality of speech generation often suffers due to challenges in cross-modal alignment. This work focuses on the LLM-centric SLMs, aiming to explore how core components affect cross-modal alignment and the quality of generated speech.
@@
We observe a mismatch in information density between speech and text representations in the speech head: one second of speech corresponds to hundreds of tokens, far exceeding the amount of text information humans typically express per second. In LLMs, multi-token prediction (MTP) is often used to accelerate inference or enrich hidden representations by enabling a single hidden state to predict multiple future tokens. To balance the information density between speech and text, we introduce MTP into SLMs, allowing each speech hidden vector to decode multiple speech tokens (for both coupled and decoupled tokenizers). We experimented with various compression ratios, ranging from one token per hidden state to twelve tokens per state. The MTP not only improved speech decoding speed by up to 12 but also significantly enhanced speech synthesis quality (the Word Error Rate decreased from 6.07 to 3.01).
@@ 2.1 Core Components
is a key component of SLMs, as it defines the atomic representations of text and speech. High-quality representations are crucial for effective cross-modal alignment, efficient generation, and preservation of acoustic details, making this component the focus of extensive research Zhan et al. (2024); Peng et al. (2024). It converts text or speech into discrete token sequences, such as for text and for speech. These sequences are ultimately combined into interleaved text-speech sequences to handle a variety of speech-language tasks, such as ASR Fathullah et al. (2024), TTS Du et al. (2024a, b), speech translation Chen et al. (2024b), and spoken question answering Zhao et al. (2024).
Speech tokenizers can be categorized into coupled and decoupled 222 For simplicity, we assume a two-way decomposition, though in practice it can be readily extended to more components, e.g., . , 333 Decoupled sequences may also be arranged in a type-wise contiguous manner within a chunk (e.g., ) rather than in an interleaved fashion. However, experiments in Appendix 8 show that interleaved ordering yields significantly better performance. designs for semantic and acoustic details . Intuitively, as illustrated in Figure 1, decoupled semantic tokens are easier to align with text tokens, while the SLM can leverage its overparameterization to handle the acoustic detail tokens specifically. In contrast, coupled tokenizers may compromise cross-modal semantic alignment.
@@ 2.2 Multi Token Prediction
Multi-token prediction (MTP) Gloeckle et al. (2024); Cai et al. (2024); Liu et al. (2024) is commonly used in LLMs to accelerate inference or enrich hidden representations (e.g., by requiring to decode not only but also future tokens like ). To compress the information density of speech representations, we introduce MTP into the SLM framework. As depicted in Figure 2, compared with Next Token Prediction (NTP), MTP enables each hidden vector to decode multiple speech tokens, whether coupled or decoupled, thus alleviates the imbalance in information density between speech and text.
We group every adjacent speech tokens into a multi-token group , which is decoded simultaneously at step , meaning all tokens in are predicted from the same hidden vector . This significantly increases the information density encoded in , making it more compatible with text tokens. The definitions of the multi-token groups / 444 Where the group consists of half semantic tokens and half acoustic detail tokens . for coupled/decoupled tokenizers subject to are as follows:
@@ 2.3 Speaker-Aware Speech Generation
Inspired by Du et al. (2024a) and Li et al. (2024), we employ a pretrained timbre extractor 555 We adopt the timbre extractor from the decoupled tokenizer FACodec Ju et al. (2024) in this work. Implementation details are provided in Appendix B. It can be smoothly replaced with other extractors if needed. to obtain the timbre representation of the speaker or role. This representation can either be a discretized token , or a continuous vector sequence , where is the sequence length. The speech-language context of the SLM is then extended as follows:
@@
3 Role-Playing Knowledge QA Task
To evaluate which SLM architecture better aligns and generalizes the inherent world knowledge embedded in LLMs, and which design better preserves speaker or character consistency, we construct a role-playing knowledge QA task. On the data side, we introduce RoleTriviaQA, a role-playing speech knowledge QA dataset. On the model training side, we propose a two-stage training framework. An overview of our proposed model is illustrated in Figure 3.
3.1 RoleTriviaQA
Inspired by the text knowledge QA— TriviaQA (2017), we introduce a role-playing speech QA extension. It requires the SLM to retain the LLM’s knowledge and then generate accurate speech consistent with a specific role. To build a high-quality role-playing knowledge QA dataset, we design a two-stage data construction pipeline: (1) Speaker selection from Genshin Impact: 10 seen speakers for both training and testing, and 5 unseen only for testing; (2) Speech synthesis via CosyVoice 2 (2024b), leveraging filtered high-quality reference speeches to synthesize the role-specific speech of the text answers.
This pipeline yields 138,384 training samples, with 1,500 in-domain and 926 out-of-domain evaluation samples. More details of RoleTriviaQA are provided in Appendix G.
Model | Codebook Size | Nq | Token Rate (TPS) | Success Rate (SR) | UTMOS | WER | SIM | |
Seen | Unseen | |||||||
Coupled Tokenizers trained with SLM | ||||||||
Encodec (2023) | 1024 | 4 | 300 | 0.38 | 2.86 | 17.00 | 0.18 | 0.10 |
Encodec (2023) | 1024 | 8 | 600 | 0.24 | 3.50 | 8.14 | 0.18 | 0.10 |
WavTokenizer (2025) | 4096 | 1 | 75 | 0.38 | 2.80 | 89.60 | 0.06 | 0.04 |
WavTokenizer-v2 (2025) | 4096 | 1 | 75 | 0.65 | 3.16 | 49.24 | 0.11 | 0.08 |
StableCodec (2024) | 15625 | 2 | 50 | 0.98 | 3.90 | 13.04 | 0.17 | 0.14 |
BigCodec (2024) | 8192 | 1 | 80 | 0.70 | 3.93 | 13.63 | 0.18 | 0.10 |
Semi-Decoupled Tokenizers trained with SLM | ||||||||
SpeechTokenizer (2024) | 1024 | 5 | 250 | 0.99 | 3.86 | 13.13 | 0.16 | 0.13 |
SpeechTokenizer (2024) | 1024 | 8 | 400 | 0.90 | 4.05 | 9.95 | 0.14 | 0.14 |
Decoupled Tokenizers trained with SLM | ||||||||
FACodec-NTP-1H | 1024 | 3 | 240 | 0.87 | 3.93 | 6.07 | 0.50 | 0.49 |
FACodec-MTP-3H | 1024 | 3 | 80 | 0.82 | 3.83 | 5.84 | 0.47 | 0.47 |
FACodec-MTP-6H | 1024 | 3 | 40 | 0.92 | 3.75 | 4.36 | 0.47 | 0.47 |
@@ 3.1 RoleTriviaQA
3.2 Training
Stage 1: Speech-Language Pretraining
To improve cross-modal alignment, we conduct a large-scale dataset Emilia-3.5. First, we filter Emilia (2024) (a large-scale, real-world speech collection with diverse emotions and styles) for samples scoring at least 3.5 on DNSMOS (2022). Next, we merge the standard LibriTTS (2019) training set. The resulting corpus comprises 2,495 hours of high-quality speech.
Moreover, to enable bidirectional alignment, we conduct both ASR and TTS training. Specifically, we adopt as the TTS data format and for ASR. The pretaining loss is formulated as:
Stage 2: Instruction Fine-Tuning
After the pretraining stage, we perform supervised fine-tuning (SFT) with RoleTriviaQA. The data format is , and the SFT loss is defined as follows:
4 Experiment
Our experiments address two critical questions:
- 1.
2.
How can SLMs effectively leverage their internal knowledge while maintaining high-quality speech synthesis? (Section 4.4)
Setups
4.1 Effect of Speech Tokenizers
4.2 Next Token vs. Multi-Token Prediction
As illustrated in Table 1, to investigate the effectiveness of MTP under decoupled architecture, we also conduct experiments comparing models trained with different compression ratios.
4.3 Speaker-Aware Speech Generation
@@
Models | Model Type | Tokenizer Type | In-Domain (RoleTriviaQA) | Out-of-Domain (Web Questions) | ||||||||
EM(%) | F1(%) | SIM | EM | F1 | SIM | |||||||
Seen | Unseen | All | Seen | Unseen | All | |||||||
Qwen2.5-0.5B-Instruct | LLM | - | 7.1 | 17.3 | - | - | - | 1.1 | 14.0 | - | - | - |
WavTokenizer-v2 | SLM | Coupled | 5.1 | 12.5 | 0.22 | 0.18 | 0.20 | 1.1 | 3.9 | 0.21 | 0.18 | 0.20 |
BigCodec | SLM | Coupled | 2.9 | 9.4 | 0.23 | 0.17 | 0.21 | 1.0 | 4.2 | 0.23 | 0.17 | 0.21 |
StableCodec | SLM | Coupled | 7.5 | 13.8 | 0.18 | 0.14 | 0.17 | 0.6 | 3.6 | 0.17 | 0.14 | 0.16 |
SpeechTokenizer | SLM | Semi-Decoupled | 3.4 | 9.1 | 0.19 | 0.14 | 0.17 | 0.3 | 2.6 | 0.20 | 0.14 | 0.18 |
@@ 4.4 Role-Playing Knowledge QA
To evaluate the capability of our decoupled SLM in preserving both knowledge and acoustic features, we compare it against the top-3 coupled and semi-decoupled SLM baselines (based on Table 1) and the LLM backbone Qwen2.5-0.5B-Instruct.
Results
Table 3 shows that the decoupled model establishes state-of-the-art performance across all metrics. First, for in-domain evaluation, our decoupled model achieves 12.0 EM and 23.8 F1, surpassing all baselines, while its 0.61 SIM score is 2.9 higher than the best coupled baseline (BigCodec). For out-of-domain evaluation, while all models degrade, ours exhibits superior robustness. This highlights that decoupling enables joint optimization of semantic and acoustic objectives, whereas semi-decoupled/coupled systems suffer from target conflicts. Second, compared with the LLM baseline, our decoupled SLM significantly outperforms on both ID and OOD evaluations. Notably, for OOD, the EM (9.0 vs 1.1) and the F1 (17.9 vs 14.0), demonstrating its effectiveness in preserving LLM knowledge while adapting to unseen queries. Particularly, for SIM generalization, our model achieves high SIM scores on both seen/unseen speakers (0.63 vs 0.58), with minimal performance gaps. This demonstrates its strong generalization in preserving speaker timbre.
Overall, our results confirm that the decoupled architecture enables cooperative improvement in knowledge QA tasks and speaker timbre, and outperforms other baselines.

### v2

# What Makes a Good Speech Tokenizer for LLM-Centric Speech Generation? A Systematic Study

@@ Abstract
Speech-language models (SLMs) offer a promising path toward unifying speech and text understanding and generation. However, challenges remain in achieving effective cross-modal alignment and high-quality speech generation. In this work, we systematically investigate the role of speech tokenizer designs in LLM-centric SLMs, augmented by speech heads and speaker modeling. We compare coupled, semi-decoupled, and fully decoupled speech tokenizers under a fair SLM framework and find that decoupled tokenization significantly improves alignment and synthesis quality. To address the information density mismatch between speech and text, we introduce multi-token prediction (MTP) into SLMs, enabling each hidden state to decode multiple speech tokens. This leads to up to 12 faster decoding and a substantial drop in word error rate (from 6.07 to 3.01). Furthermore, we propose a speaker-aware generation paradigm and introduce RoleTriviaQA, a large-scale role-playing knowledge QA benchmark with diverse speaker identities. Experiments demonstrate that our methods enhance both knowledge understanding and speaker consistency.
Code
https://github.com/cnxupupup/SLM-Decoupled-MTP
Demo
https://cnxupupup.github.io/SLM-Decoupled-MTP-Demo
@@ 1 Introduction
In recent years, large language models (LLMs) have achieved remarkable progress in text understanding and generation (Grattafiori et al. 2024; Qwen et al. 2025), fueling rapid advancements in multimodal models such as vision-language models (Liu et al. 2023), vision-language-action models (Kim et al. 2024), and speech-language models (SLMs) (Chu et al. 2023; Li et al. 2025). Among them, SLMs are particularly valuable for applications like natural and fluent human-computer dialogue (Défossez et al. 2024; Xu et al. 2025) and personalized speech generation (Du et al. 2024a, c). The core challenge lies in generalizing the language understanding and generation capabilities of LLMs to incorporate and integrate speech understanding and generation.
Current approaches for connecting text and speech in SLMs can be broadly categorized into two perspectives. Speech-centric models (Du et al. 2024a, c; Anastassiou et al. 2024) are built upon mature speech generation architectures, with LLMs introduced as conditional context. However, while this approach achieves high-quality speech synthesis, it struggles to fully leverage LLMs’ rich world knowledge and powerful capabilities. LLM-centric models (Défossez et al. 2024; Zhang et al. 2023; Zhan et al. 2024) take LLMs as the backbone and extend them with speech interaction. In theory, it fully utilizes LLMs’ strengths (e.g., intent understanding, instruction following, planning, and decision-making). However, the quality of speech generation often suffers due to challenges in cross-modal alignment. This work focuses on the LLM-centric SLMs, aiming to explore how core components affect cross-modal alignment and the quality of generated speech.
@@
We observe a mismatch in information density between speech and text representations in the speech head: one second of speech corresponds to hundreds of tokens, far exceeding the amount of text information humans typically express per second. In LLMs, multi-token prediction (MTP) (Gloeckle et al. 2024; Cai et al. 2024; DeepSeek-AI et al. 2024) is often used to accelerate inference or enrich hidden representations by enabling a single hidden state to predict multiple future tokens. To balance the information density between speech and text, we introduce MTP into SLMs, allowing each speech hidden vector to decode multiple speech tokens (for both coupled and decoupled tokenizers). We experimented with various compression ratios, ranging from one token per hidden state to twelve tokens per state. The MTP not only improved speech decoding speed by up to 12 but also significantly enhanced speech synthesis quality (the Word Error Rate decreased from 6.07 to 3.01).
@@ 2.1 Core Components
is a key component of SLMs, as it defines the atomic representations of text and speech. High-quality representations are crucial for effective cross-modal alignment, efficient generation, and preservation of acoustic details, making this component the focus of extensive research (Zhan et al. 2024; Peng et al. 2024). It converts text or speech into discrete token sequences, such as for text and for speech. These sequences are ultimately combined into interleaved text-speech sequences to handle a variety of speech-language tasks, such as ASR (Fathullah et al. 2024), TTS (Du et al. 2024a, b), speech translation (Chen et al. 2024b), and spoken question answering (Zhao et al. 2024).
Speech tokenizers can be categorized into coupled and decoupled 111 For simplicity, we assume a two-way decomposition, though in practice it can be readily extended to more components, e.g., . , 222 Decoupled sequences may also be arranged in a type-wise contiguous manner within a chunk (e.g., ) rather than in an interleaved fashion. However, experiments in Section F.3 show that interleaved ordering yields significantly better performance. designs for semantic and acoustic details . Intuitively, as illustrated in Figure 1, decoupled semantic tokens are easier to align with text tokens, while the SLM can leverage its overparameterization to handle the acoustic detail tokens specifically. In contrast, coupled tokenizers may compromise cross-modal semantic alignment.
@@ 2.2 Multi Token Prediction
Multi-token prediction (MTP) (Gloeckle et al. 2024; Cai et al. 2024; Liu et al. 2024) is commonly used in LLMs to accelerate inference or enrich hidden representations (e.g., by requiring to decode not only but also future tokens like ). To compress the information density of speech representations, we introduce MTP into the SLM framework. As depicted in Figure 2, compared with Next Token Prediction (NTP), MTP enables each hidden vector to decode multiple speech tokens, whether coupled or decoupled, thus alleviates the imbalance in information density between speech and text.
We group every adjacent speech tokens into a multi-token group , which is decoded simultaneously at step , meaning all tokens in are predicted from the same hidden vector . This significantly increases the information density encoded in , making it more compatible with text tokens. The definitions of the multi-token groups / 333 Where the group consists of half semantic tokens and half acoustic detail tokens . for coupled/decoupled tokenizers subject to are as follows:
@@ 2.3 Speaker-Aware Speech Generation
Inspired by Du et al. (2024a) and Li et al. (2024), we employ a pretrained timbre extractor 444 We adopt the timbre extractor from the decoupled tokenizer FACodec (Ju et al. 2024) in this work. Implementation details are provided in Appendix B. It can be smoothly replaced with other extractors if needed. to obtain the timbre representation of the speaker or role. This representation can either be a discretized token , or a continuous vector sequence , where is the sequence length. The speech-language context of the SLM is then extended as follows:
@@
Model |
Nq |
UTMOS | WER | SIM | |||||||
Seen | Unseen | |||||||||||||
Coupled Tokenizers trained with SLM | ||||||||||||||
Encodec (2023) | 1024 | 4 | 300 | 0.38 | 2.86 | 17.00 | 0.18 | 0.10 | ||||||
Encodec (2023) | 1024 | 8 | 600 | 0.24 | 3.50 | 8.14 | 0.18 | 0.10 | ||||||
WavTokenizer (2025) | 4096 | 1 | 75 | 0.38 | 2.80 | 89.60 | 0.06 | 0.04 | ||||||
WavTokenizer-v2 (2025) | 4096 | 1 | 75 | 0.65 | 3.16 | 49.24 | 0.11 | 0.08 | ||||||
StableCodec (2024) | 15625 | 2 | 50 | 0.98 | 3.90 | 13.04 | 0.17 | 0.14 | ||||||
BigCodec (2024) | 8192 | 1 | 80 | 0.70 | 3.93 | 13.63 | 0.18 | 0.10 | ||||||
Semi-Decoupled Tokenizers trained with SLM | ||||||||||||||
SpeechTokenizer (2024) | 1024 | 5 | 250 | 0.99 | 3.86 | 13.13 | 0.16 | 0.13 | ||||||
SpeechTokenizer (2024) | 1024 | 8 | 400 | 0.90 | 4.05 | 9.95 | 0.14 | 0.14 | ||||||
Decoupled Tokenizers trained with SLM | ||||||||||||||
FACodec-NTP-1H | 1024 | 3 | 240 | 0.87 | 3.93 | 6.07 | 0.50 | 0.49 | ||||||
FACodec-MTP-3H | 1024 | 3 | 80 | 0.82 | 3.83 | 5.84 | 0.47 | 0.47 | ||||||
FACodec-MTP-6H | 1024 | 3 | 40 | 0.92 | 3.75 | 4.36 | 0.47 | 0.47 | ||||||
@@
3 Role-Playing Knowledge QA Task
Current SLMs excel at daily conversations (Défossez et al. 2024), but their performance on knowledge-intensive tasks (e.g., QA) and voice characteristics (e.g., role-playing) lacks systematic evaluation. To address this problem, we introduce the Role-Playing Knowledge QA task, assessing SLM knowledge retention from its backbone LLM and similarity to reference speech in voice characteristics.
However, a critical gap exists: no open-source dataset assesses both correctness and timbre similarity in speech knowledge QA. Other models featuring speech QA (Défossez et al. 2024) trained and evaluated on a large-scale data which has not been released yet. To bridge this gap and stimulate SLMs’ intrinsic knowledge with substantial training (Zhang et al. 2023), we propose RoleTriviaQA, an open-source dataset for benchmarking SLMs on role embodiment and knowledgeable spoken responses. Similar with (Zhang et al. 2023), we adopt the data format (see Figure 3), which enables essential text-speech knowledge alignment for high-quality role-playing.
3.1 Data Construction
We created the RoleTriviaQA dataset by selecting 15 distinct voices from Genshin Impact555https://genshin.hoyoverse.com/en roles (10 seen: 5 Male/5 Female for training and evaluation; 5 unseen: 2 Male/3 Female for evaluation only). Then we synthesize speech for TriviaQA (Joshi et al. 2017) QA-pairs using CosyVoice 2 (Du et al. 2024b) after filtering high-quality reference speeches. This pipeline yields 138,384 training samples and a test set with 1,500 in-domain (TriviaQA) and 926 out-of-domain (OpenAudioBench (Li et al. 2025)) samples.Further details of RoleTriviaQA are in Appendix H.
3.2 Training Strategy
Stage 1: Cross-Modal Alignment Pretraining
To improve cross-modal alignment, we conduct a large-scale dataset Emilia-3.5. First, we filter Emilia (2024) (a large-scale, real-world speech collection with diverse emotions and styles) for samples scoring at least 3.5 on DNSMOS (2022). Next, we merge the standard LibriTTS (2019) training set. The resulting corpus comprises 2,495 hours of high-quality speech. Moreover, To enable bidirectional alignment of text and speech modalities, we conduct pretrain on both TTS and ASR tasks. Specifically, we adopt as the TTS data format and for ASR. The pretaining loss is formulated as:
Stage 2: Knowledge-Role Joint Fine-Tuning
After the pretraining stage, we perform supervised fine-tuning (SFT) with RoleTriviaQA to enable the SLM to capture the voice features of each role along with the knowledge-intensive data.
The SFT loss is defined as follows:
4 Experiment
Models |
In-Domain (RoleTriviaQA) | Out-of-Domain (Web Questions) | ||||||||||||
EM(%) | F1(%) | SIM | EM | F1 | SIM | |||||||||||
Seen | Unseen | All | Seen | Unseen | All | |||||||||||
Qwen2.5-0.5B-Instruct | LLM | - | 7.1 | 17.3 | - | - | - | 1.1 | 14.0 | - | - | - | ||||
WavTokenizer-v2 | SLM | Coupled | 5.1 | 12.5 | 0.22 | 0.18 | 0.20 | 1.1 | 3.9 | 0.21 | 0.18 | 0.20 | ||||
BigCodec | SLM | Coupled | 2.9 | 9.4 | 0.23 | 0.17 | 0.21 | 1.0 | 4.2 | 0.23 | 0.17 | 0.21 | ||||
StableCodec | SLM | Coupled | 7.5 | 13.8 | 0.18 | 0.14 | 0.17 | 0.6 | 3.6 | 0.17 | 0.14 | 0.16 | ||||
SpeechTokenizer | SLM | Semi-Decoupled | 3.4 | 9.1 | 0.19 | 0.14 | 0.17 | 0.3 | 2.6 | 0.20 | 0.14 | 0.18 | ||||
@@ 4 Experiment
Our experiments address two critical questions:
- 1.
2.
How can SLMs effectively leverage their intrinsic knowledge while maintaining high-quality speech synthesis? (Section 4.4)
Setups
4.1 Effect of Speech Tokenizers
4.2 Next Token vs. Multi-Token Prediction
As shown in Table 1, to investigate the effectiveness of MTP under decoupled architecture, we also conduct experiments comparing models trained with different compression ratios.
4.3 Speaker-Aware Speech Generation
@@ 4.4 Role-Playing Knowledge QA
We evaluate Role-Playing Knowledge QA within two aspects: 1) Benchmark of the decoupled SLM against top-3 coupled/semi-decoupled SLMs (From Table 1) and the LLM backbone(Qwen2.5-0.5B-Instruct); 2) Compare the decoupled SLM against diverse SLM baselines.
Results on Different Tokenizers
As shown in Table 3, SLM trained with decoupled tokenizer achieves SOTA performance across all metrics. For in-domain evaluation, it outperforms all baselines with 12.0 EM and 23.8 F1, while its 0.61 SIM score is 2.9× higher than the best coupled baseline (BigCodec). Though all models degrade on OOD queries, the decoupled model shows greater robustness, confirming that decoupling resolves target conflicts in semi-decoupled/coupled systems by jointly optimizing semantic and acoustic objectives.
Compared to the LLM baseline, the decoupled SLM significantly outperforms in both ID/OOD settings (OOD EM: 9.0 vs 1.1; F1: 17.9 vs 14.0), effectively preserving LLM knowledge while adapting to unseen queries. It also generalizes well in speaker timbre, achieving high SIM scores for both seen/unseen speakers (0.63 vs 0.58) with minimal gaps.
Overall, the decoupled architecture thus enables cooperative improvement in knowledge QA and speaker timbre, surpassing all baselines.
Results on Different SLMs
Comparisons with other SLM baselines also confirm that decoupled tokenizers help not only improve the speech quality but preserve LLM knowledge. Detailed results are provided in Appendix E.